OpenStudy (anonymous):

solve for x using quadratic function 2x/5=(x+1)/3x+1

OpenStudy (anonymous):

cross multiply to get 2x(3x+1) = 5(x+1) Distribute to get 6x^2 + 2x = 5x + 5 Then move everything to the right hand side to get 6x^2 - 3x - 5 = 0 Then solve the quadratic.

OpenStudy (anonymous):

so subtract 6x from 3x?

OpenStudy (anonymous):

no, because they are not like terms. do you know how to solve a quadratic equation by factoring or by using the quadratic formula?

OpenStudy (anonymous):

not really and i just need to see the steps then i will be ok

OpenStudy (anonymous):

hope you understood what the steps did so far. if not, feel free to ask. to continue - we need to use the quadratic formula - which is x=\[(-b \pm \sqrt{b ^{2}-4ac})/2a\] and the a is the coeff of x^2, b is coeff of x and c is the constant from the equation we have,

OpenStudy (anonymous):

so, in your problem, a = 6, b = -3, and c = -5. just plug them into the formula to find the two possible values of x.

OpenStudy (anonymous):

how did you get rid of the ^2

OpenStudy (anonymous):

when you solve a quadratic, you are finding the value of x. there are two ways to do this. one is to factor, and the other is to use the formula mentioned above. either way, gives you a value of x.

OpenStudy (anonymous):

did you get an answer?

OpenStudy (anonymous):

ok im doing it now

OpenStudy (anonymous):

i got 3+ square root of 129/ 12

OpenStudy (anonymous):

hold up thats not right

OpenStudy (anonymous):

great. don't forget that it could also be (3-sqrt129)/12. the sign is the plus-or-minus sign. almost every time a quadratic is solved, you get two possible answers for x.

OpenStudy (anonymous):

nvm how do you know if its right?

OpenStudy (anonymous):

generally, you can plug into the original and check your work. in this case, it would be difficult. can you check the answers? have you done problems like these where you get weird answers? my strategy, if all else fails, is to make sure that the problem i started out with is correct (no sign errors in the original.)

OpenStudy (anonymous):

thanks for the medal, by the way!

OpenStudy (anonymous):

I want to no if this is right? x/5=3x+1/2 15x+5=2x is this one not possible?

OpenStudy (anonymous):

is (3x + 1) all being divided by 2, or just the 1?

OpenStudy (anonymous):

all of it

OpenStudy (anonymous):

then you are correct, you will get 15x + 5 = 2x. this is not a quadratic, because there is no x^2. Just use algebra to get all the x-terms to one side, getting 5 = -13x then divide by -13 to get x = -5/13

OpenStudy (anonymous):

ok thats what i got thanks

OpenStudy (anonymous):

you are welcome.